a ball ‘A’ of mass m falls to the surface of earth from infinity.Another ball ‘B’of mass 2m falls to the earth from the height equal to six times radius of earth.What would be the ratios of velocities of ‘A’ to’B’ on reaching the earth

Grade:11

3 Answers

DR STRANGE
7 years ago
HI AVNEET,
 YOU CAN SOLVE IT BY CONCEPT OF CONSERVTION OF ENERGY,
CHANGE IN POTENTIAL ENERGY IS KINETIC ENERGY GAINED IN THIS CASE.
FOR BALL “B”,
K.E.=P2-P1=m_{b}gR-m_{b} gR(\frac{R^2}{R^{2}+H^{2}}) = m_{b}gR\left ( \frac{H^{2}}{R^{2}+H^{2}} \right )= \frac{1}{2}m_{b}v_{b}^{2}
 
 
FOR BALL “A”,K.E.= P2-P1=P2   AS AT INFINITY POTENTIAL ENERGY IS ZERO.
K.E.= \frac{1}{2}m_{a}v_{a}^{2}=m_{a}gR
SO JUST TAKE THE RATIO AND SQUARE ROOT ON BOTH SIDES PUT THE VALUES ma, mb , AND H=6R 
YOU WILL GET THE ANSWER
PLEASE APPROVE THE ANSWER
ALL THE BEST AVNEET :)
 
 
DR STRANGE
7 years ago
SORRY IN FIRST EQUATION 
P1= m_{b}gH\left ( \frac{R^{2}}{R^{2}+H^{2}} \right ),
PERFORM THE MATHS AGAIN
ASKIITIANS SHOULD GIVE AN EDIT OPTION FOR THE ANSWERS
 
SORRY FOR THE MISTAKE I HOPE YOU WILL GET IT
DR STRANGE
7 years ago
 
IGNORE WHOLE ANSWER ITS INCORRECT! 
IGNORE WHOLE ANSWER ITS INCORRECT! 
IGNORE WHOLE ANSWER ITS INCORRECT!